# EmT: A Novel Transformer for Generalized Cross-subject EEG Emotion Recognition

### v1

@@ V-F Visualization
Figure 5 illustrates the learned connectivity of brain regions on SEED. Two learnable adjacency matrices reveal different connectivity patterns during the emotional cognitive process. Figure 5 clearly shows a difference in the learned connectivities under emotion stimulation. In Figure 5 (a), the strongest learned connections are F6-POz, P2-Fz, F7-O1, PO5-Pz, T8-P7, and FC4-FPz. These connections indicate relationships among the frontal, parietal, and temporal areas, which are known to be closely related to mental attention [37]. In Figure 5 (b), the important connections learned by EmT are FC4-P7, FT8-C5, Cz-Fz, T8-CP4, CP6-FP2, F2-TP8, and O1-P2. These connections include interactions among the frontal, temporal, and parietal areas, which are known to be related to emotions [38, 39, 40]. Additionally, some interactions involve the occipital and parietal areas, which are associated with visual processes. This is expected as videos were used as stimuli in the data collection experiments [22]. The two distinct patterns in Figure 5 reflect EmT’s ability to capture multiple cognitive connectivities for the given classification task.
@@ VI Conclusion
In this paper, we propose a graph-transformer-based model, EmT, for cross-subject EEG emotion recognition. RMPG is proposed to learn multiple connection patterns for different brain cognitive processes under emotional stimulation. TCT is designed to learn temporal contextual information from the temporal EEG graphs. Subject-independent classification and regression tasks are conducted to evaluate EmT and relevant baseline methods. Results on three bench-marking datasets demonstrate EmT shows improvements over the compared baselines. Extensive analyses are performed to interpret the results and neurophysiologically meaningful features learned. Despite the improved emotion recognition performance, the learned temporal contextual information is difficult to identify. Proper experiments should be designed to pinpoint this temporal contextual information in the future. Additionally, in this paper, spatial connection information is learned using GCNs. Another useful tool for capturing connection patterns in EEG is the Riemannian manifold. Therefore, exploring how to capture contextual information using functional connection features and Riemannian manifolds would be an interesting direction for future research.

### v2

@@ V-H1 The learned connections
Fig. 6 illustrates the learned connectivity of brain regions on SEED. Two learnable adjacency matrices reveal different connectivity patterns during the emotional cognitive process. Fig. 6 clearly shows a difference in the learned connectivities under emotion stimulation. In Fig.  6 (a), the strongest learned connections are F6-POz, P2-Fz, F7-O1, PO5-Pz, T8-P7, and FC4-FPz. These connections indicate relationships among the frontal, parietal, and temporal areas, which are known to be closely related to mental attention [38]. In Fig. 6 (b), the important connections learned by EmT are FC4-P7, FT8-C5, Cz-Fz, T8-CP4, CP6-FP2, F2-TP8, and O1-P2. These connections include interactions among the frontal, temporal, and parietal areas, which are known to be related to emotions [39, 40, 41]. Additionally, some interactions involve the occipital and parietal areas, which are associated with visual processes. This is expected as videos were used as stimuli in the data collection experiments [22]. The two distinct patterns in Fig. 6 reflect EmT’s ability to capture multiple cognitive connectivities for the given classification task.
V-H2 The learned temporal contextual information
We visualized the hidden features before and after the TCT blocks to examine 1) the nature of temporal contextual information and 2) how TCT uses it for classification and regression tasks. For classification, we visualized the averaged features of samples within the same class for each subject, with a representative subject in FACED used to demonstrate the learned temporal contextual information. For regression, we visualized individual sample features, as each has a unique ground truth, and a representative sample is used for illustration purposes.
The visualization results for the classification task are shown in Fig. 8. Here, $d$ represents the dimension of hidden features, and $t$ denotes time. $S_{T}\in\mathbb{R}^{seq\times d_{g}}$ and $S_{cls}\in\mathbb{R}^{seq\times d_{head}}$ represent the learned features before and after the TCT blocks, respectively, with superscripts indicating class labels. As seen in Fig. 8, the features change over time before the TCT-Clas blocks, particularly for class 0 (low valence stimuli). After the TCT blocks, the activations become more consistent, likely due to self-attention focusing on parts strongly correlated with the emotional state and the SAT layer smoothing fluctuations by aggregating nearby temporal information. For regression, the visualization (Fig. 9) also shows temporal changes in the feature space, indicating the presence of temporal contextual information. Unlike classification, the regression features are not simply smoothed. Instead, TCT-Regr, using RNNs as token mixers, combines evolving features across different dimensions, resulting in a more complex representation that is well-suited for continuous emotional state prediction. The RNN’s ability to capture sequential dependencies allows it to focus on key changes necessary for continuous prediction.
In summary, the changes in features before the TCT blocks confirm the presence of temporal contextual information reflecting the dynamic nature of emotional states. The TCT blocks handle tasks differently: in classification, features are smoothed to enhance separability, while in regression, RNN token mixers preserve temporal variations, enabling continuous emotional predictions.
(a) SEED (62 Channels)
(b) FACED (32 Channels)
Figure 7: Comparison of model performance, parameters, and FLOPS (M). The model size is represented by the area of the circle. TSception and Conformer are excluded due to much higher FLOPS than others. Our EmT-B achieves a promising balance between performance and computational complexity.
V-I Limitations and Future Work
Although our model performs better overall across datasets, there are some limitations. First, the model’s size and computational complexity can be further optimized. Additionally, while our model improves emotion recognition, the temporal contextual information learned remains difficult to interpret. Future work should design experiments to pinpoint this information. Exploring alternative methods, such as using the Riemannian manifold to capture EEG connection patterns, would also be a promising research direction.
(a) Low valance
(b) High valance
Figure 8: Visualization of the learned temporal features before and after TCT blocks for the classification task. $d$ represents the dimension of hidden features, and $t$ denotes time. $S_{T}\in\mathbb{R}^{seq\times d_{g}}$ and $S_{cls}\in\mathbb{R}^{seq\times d_{head}}$ represent the learned features before and after the TCT blocks, respectively, with superscripts indicating class labels.
Figure 9: Visualization of the learned temporal features before and after TCT blocks for the regression task. $d$ represents the dimension of hidden features, and $t$ denotes time. $S_{T}\in\mathbb{R}^{seq\times d_{g}}$ and $S_{reg}\in\mathbb{R}^{seq\times d_{head}}$ represent the learned features before and after the TCT blocks.
@@ VI Conclusion
In this paper, we propose a graph-transformer-based model, EmT, for cross-subject EEG emotion recognition. RMPG is proposed to learn multiple connection patterns for different brain cognitive processes under emotional stimulation. TCT is designed to learn temporal contextual information from the temporal EEG graphs. Subject-independent classification and regression tasks are conducted to evaluate EmT and relevant baseline methods. Results on three bench-marking datasets demonstrate EmT shows improvements over the compared baselines. Our approach is inspired by neuropsychological knowledge and offers a new perspective for improving decoding performance by learning multiple cognition-related graph connections and capturing temporal contextual information. These improvements are supported by enhanced performance and visualizations of the learned connections and temporal features. Additionally, the visualization of the temporal features provides insights into how different token mixers function in classification and regression tasks.